a scale drawing has a scake of 1 in.:10 ft. what is the distance on teh drawing for an actual distance of 20ft?of 45 ft?
Answer by checkley71(8403)

The ratio of 1 in. = 10ft. 1/10 = 2/20 or 20 ft. = 2in. 1/10 = 4.5/45 or
45 ft. = 4.5 in.